# Ratio asymptotics and zero density for orthogonal polynomials with varying Verblunsky coefficients

### v1

@@ 4.3. Asymptotically periodic variable Verblunsky coefficients
Here we establish a generalization of Theorem 10 to asymptotically periodic setting. Both approaches presented earlier, the one based on the limiting logarithmic potential and the moment method, extend naturally to this extended framework.
@@
$\sigma_{t}=\frac{1}{t}\int_{0}^{t}\nu_{\Delta_{s}}{\rm d}s.$ (4.26)
@@
$\mathop{\mathrm{supp}}\nolimits\nu_{\Delta}=\left\{e^{{\rm i}\theta}\mid\theta\in[\theta^{(+)},\theta^{(-)}]\cup[-\theta^{(-)},-\theta^{(+)}]\right\},$
@@
$\frac{{\rm d}\nu_{\Delta}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi}\frac{|\sin\theta|}{\sqrt{\left(\cos\theta^{(+)}-\cos\theta\right)\left(\cos\theta-\cos\theta^{(-)}\right)}},$
for $\pm\theta\in[\theta^{(+)},\theta^{(-)}]$; see [simon_opuc2, Ex. 11.1.5]. Then similarly as in Remark 11, we can deduce the 2-periodic generalization of the formula for the density (4.4) of the average measure (4.24) at the point $e^{{\rm i}\theta}$ which reads
@@ 5. Asymptotic zero density for variable OPUC
In this section, we extend the ideas of Theorem 10 to the setting of zeros of OPUC. Recall from Section 2.7 that, even in the non-varying case, the behavior of OPUC zeros is more intricate than that of POPUC. In particular, when $\alpha_{n}\to 0$, the rate of convergence becomes important. More concretely, we prove a variable generalization of the fact that the asymptotic distribution of zeros of OPUC is given by the normalized Lebesgue measure on $\partial\mathbb{D}$} if $\alpha_{n}\to 0$ with $|\alpha_{n}|^{1/n}\to 1$. On the other, if the latter condition is altered to $|\alpha_{n}|^{1/n}\to A\in[0,1)$ as in the Mhaskar–Saff theorem, we do not have a variable generalization. This remains an open problem (see Remark 22 below).
@@ 5.1. Balayage
$\int_{\overline{{\mathbb{D}}}}z^{k}{\rm d}\mu(z)=\int_{0}^{2\pi}e^{ik\theta}{\rm d}\mathcal{P}(\mu)(e^{{\rm i}\theta}),\quad\forall k\in\mathbb{N}_{0}.$ (5.1)
@@
$\int_{{\partial\mathbb{D}}}f(z){\rm d}\mathcal{P}(\mu)(z)=\int_{\overline{{\mathbb{D}}}}\mathcal{P}_{*}f(z){\rm d}\mu(z),\quad\forall f\in C({\partial\mathbb{D}}),$
where
$\mathcal{P}_{*}f(re^{{\rm i}\theta})=\begin{cases}\frac{1}{2\pi}\int_{0}^{2\pi}P_{r}(\theta-t)f(e^{{\rm i}t}){\rm d}t,&\;\mbox{ if }r<1,\\[3.0pt] f(e^{{\rm i}\theta}),&\;\mbox{ if }r=1,\end{cases}$
@@
$P_{r}(t)=\sum_{n=-\infty}^{\infty}r^{|n|}e^{{\rm i}nt}=\frac{1-r^{2}}{1-2r\cos t+r^{2}}$
@@
Denote by $\nu_{n,N}$ the zero-counting measure of $\Phi_{n,N}$, i.e.,
$\nu_{n,N}=\frac{1}{n}\sum_{k=1}^{n}\delta_{z_{k}},$
where $z_{k}=z_{k,n,N}$, $k\in\{1,\dots,n\}$, are zeros of $\Phi_{n,N}$, counted with their multiplicities.
@@
Since a negative moment of a positive measure supported on $\partial\mathbb{D}$} is the complex-conjugate of the respective positive moment and Laurent polynomials are dense in $C({\partial\mathbb{D}})$, it is sufficient to show the convergence of the positive moments of the measures in (5.2). By (5.1), the $k$-th moment of $\mathcal{P}(\nu_{n,N})$ is
$\int_{0}^{2\pi}e^{{\rm i}k\theta}{\rm d}\mathcal{P}(\nu_{n,N})\big(e^{{\rm i}\theta}\big)=\int_{\overline{{\mathbb{D}}}}z^{k}{\rm d}\nu_{n,N}(z)=\frac{1}{n}\mathop{\mathrm{Tr}}\nolimits\left(\mathcal{C}_{n,N}\right)^{k}=\frac{1}{n}\mathop{\mathrm{Tr}}\nolimits\left(\mathcal{C}^{(\beta)}_{n,N}\right)^{k}+o(1),$
@@ 5.2. Asymptotic zero distribution of OPUC with variable Verblunsky coefficients
$\tilde{\nu}_{n}=\frac{1}{n}\sum_{j=1}^{n}{\delta}_{e^{{\rm i}\theta_{j}}},$
where we write $z_{j}=r_{j}e^{{\rm i}\theta_{j}}$ for some $r_{j}\geq 0$ and $\theta_{j}\in(-\pi,\pi]$; if $r_{j}=0$ then $\theta_{j}$ can be taken arbitrarily. In the first part of the proof, we show that
$\operatorname*{w-lim\ }_{n\to\infty}(\nu_{n}-\tilde{\nu}_{n})=0,$ (5.4)
@@
$\operatorname*{w-lim\ }_{n\to\infty}(\tilde{\nu}_{n}-\mathcal{P}(\nu_{n}))=0.$ (5.5)
Then (5.4) and (5.5) immediately imply the statement.
@@
$\frac{1}{n}\sum_{j=1}^{n}\left|f(z_{j})-f(e^{{\rm i}\theta_{j}})\right|=\frac{1}{n}\sum_{|z_{j}|\leq 1-\delta}\left|f(z_{j})-f(e^{{\rm i}\theta_{j}})\right|+\frac{1}{n}\sum_{|z_{j}|>1-\delta}\left|f(z_{j})-f(e^{{\rm i}\theta_{j}})\right|.$ (5.7)
@@
$\frac{1}{n}\sum_{|z_{j}|\leq 1-\delta}\left|f(z_{j})-f(e^{{\rm i}\theta_{j}})\right|\leq\frac{2\|f\|_{\overline{{\mathbb{D}}}}}{n}\,\sharp\{j\mid|z_{j}|\leq 1-\delta\},$
@@
$\lim_{n\to\infty}\frac{1}{n}\sum_{|z_{j}|\leq 1-\delta}\left|f(z_{j})-f(e^{{\rm i}\theta_{j}})\right|=0$
@@
$\frac{1}{n}\sum_{|z_{j}|>1-\delta}\left|f(z_{j})-f(e^{{\rm i}\theta_{j}})\right|\leq\frac{\epsilon}{n}\,\sharp\{j\mid|z_{j}|>1-\delta\}\leq\epsilon$
@@
$\left|z_{j}-e^{{\rm i}\theta_{j}}\right|=1-|z_{j}|<\delta.$
@@
$\lim_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}\left|f(z_{j})-f(e^{{\rm i}\theta_{j}})\right|=0,$
@@
2) The verification of (5.5): Both measures $\tilde{\nu}_{n}$ and $\mathcal{P}(\nu_{n})$ are positive measures supported in $\partial\mathbb{D}$}. Therefore, in order to obtain (5.5), it is sufficient to show that
$\lim_{n\to\infty}\int_{0}^{2\pi}e^{{\rm i}kt}{\rm d}\left(\tilde{\nu}_{n}-\mathcal{P}(\nu_{n})\right)(e^{{\rm i}t})=0,\quad\forall k\in\mathbb{N}_{0}.$ (5.8)
Let $k\in\mathbb{N}_{0}$ be fixed. Then one can make use of (5.1) to obtain
$\int_{0}^{2\pi}e^{{\rm i}kt}\,{\rm d}\!\left(\tilde{\nu}_{n}-\mathcal{P}(\nu_{n})\right)(e^{{\rm i}t})=\int_{0}^{2\pi}e^{{\rm i}kt}{\rm d}\tilde{\nu}_{n}(e^{{\rm i}t})-\int_{\overline{{\mathbb{D}}}}z^{k}{\rm d}\nu_{n}(z)=\frac{1}{n}\sum_{j=1}^{n}\left(e^{{\rm i}k\theta_{j}}-z_{j}^{k}\right).$ (5.9)
@@
$\frac{1}{n}\sum_{j=1}^{n}\left|e^{{\rm i}k\theta_{j}}-z_{j}^{k}\right|$ $=\frac{1}{n}\sum_{|z_{j}|\leq 1-\delta}\left|e^{{\rm i}k\theta_{j}}-z_{j}^{k}\right|+\frac{1}{n}\sum_{|z_{j}|>1-\delta}\left|e^{{\rm i}k\theta_{j}}-z_{j}^{k}\right|$
@@
$\operatorname*{w-lim\ }_{n/N\to t}\nu_{n,N}=\frac{1}{t}\int_{0}^{t}\nu_{|\alpha(s)|}{\rm d}s.$ (5.12)
@@
The idea of the proof relies on showing that the property (5.3) holds in the variable setting, i.e, the roots $z_{k,n,N}$, $k=1,\dots,n$, of $\Phi_{n,N}$ are approaching the unit circle up to $o(n)$ exceptions as $n/N\to t$.
@@
$\frac{1}{n_{j}}\log|\alpha_{n_{j}-1,N_{j}}|=\frac{1}{n_{j}}\sum_{k=1}^{n_{j}}\log|z_{k,n_{j},N_{j}}|\leq\frac{\log(1-\delta)}{n_{j}}\,\sharp\{k\mid|z_{k,n_{j},N_{j}}|\leq 1-\delta\},$ (5.13)
@@
It would be very interesting to establish the variable version of the Mhaskar–Saff theorem, that is, to determine the limiting zero density of OPUC when the limit in (5.11) is no longer assumed to be $1$. Addressing this problem likely requires a different approach. We leave this question as an open problem for now.
@@ 6. Examples
As an illustration, we provide examples of POPUC and OPUC with variable Verblunsky coefficients whose asymptotic zero distribution is determined by (4.2) and can be computed in an even more explicit form. This concerns particularly cases when $|\alpha|$ is monotonic in $(0,t)$.
@@
    If $f$ is strictly increasing in $(0,t)$, then $\sigma_{t}$ is the absolutely continuous measure with $\mathop{\mathrm{supp}}\nolimits\sigma_{t}=\{e^{{\rm i}\theta}\mid\theta\in[\theta_{f(0)},2\pi-\theta_{f(0)}]\}$ and
    $\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\int_{0}^{\min(t,f^{-1}(\sin(\theta/2))}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-f^{{2}}(s)}}{\rm d}s,$ (6.1)
@@
    If $f$ is strictly decreasing in $(0,t)$, then $\sigma_{t}$ is the absolutely continuous measure with $\mathop{\mathrm{supp}}\nolimits\sigma_{t}=\{e^{{\rm i}\theta}\mid[\theta_{f(t)},2\pi-\theta_{f(t)}]\}$ and
    $\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\int_{\max(0,f^{-1}(\sin(\theta/2))}^{t}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-f^{{2}}(s)}}{\rm d}s,$ (6.2)
@@
Here $\sigma_{t}$ is the average measure (4.3) and $\theta_{a}$ is defined by (2.20).
@@
Recall that $\mathop{\mathrm{supp}}\nolimits\nu_{f(s)}=\{e^{{\rm i}\theta}\mid\theta\in[\theta_{f(s)},2\pi-\theta_{f(s)}]\}$ and therefore, assuming $f$ is increasing, it follows readily from (4.3) that $\mathop{\mathrm{supp}}\nolimits\sigma_{t}=\{e^{{\rm i}\theta}\mid\theta\in[\theta_{f(0)},2\pi-\theta_{f(0)}]\}$.
@@
$F_{\sigma_{t}}(\xi)=\sigma_{t}\left(\left\{e^{{\rm i}\theta}\mid\theta\in\left[\theta_{f(0)},\xi\right]\right\}\right)=\frac{1}{t}\int_{0}^{t}\nu_{f(s)}\left(\left[\theta_{f(0)},\xi\right]\right){\rm d}s.$
@@
Suppose $f(s)=s^{\omega}$ and $0<t\leq 1$, where $\omega>0$ is a parameter. This situation corresponds to the choice $\alpha_{n,N}=\left(n/N\right)^{\omega}$ for $0\leq n\leq N$, for example. Clearly, $f$ is strictly increasing and, by Proposition (23), $\mathop{\mathrm{supp}}\nolimits\sigma_{t}={\partial\mathbb{D}}$.
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\int_{0}^{t}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-s^{2\omega}}}{\rm d}s\\ =\frac{1}{2\pi t}(\sin\left(\theta/2\right))^{1/\omega}\int_{0}^{t/(\sin(\theta/2))^{1/\omega}}\frac{{\rm d}u}{\sqrt{1-u^{2\omega}}},$ (6.3)
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi}\,_{2}F_{1}\left(\frac{1}{2},\frac{1}{2\omega};\frac{2\omega+1}{2\omega};\frac{t^{2\omega}}{\sin^{2}(\theta/2)}\right).$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\left(\sin\left(\theta/2\right)\right)^{1/\omega}\,_{2}F_{1}\left(\frac{1}{2},\frac{1}{2\omega};\frac{2\omega+1}{2\omega};1\right)\\ =\frac{1}{2\sqrt{\pi}t}\frac{\Gamma\left((2\omega+1)/(2\omega)\right)}{\Gamma\left((\omega+1)/(2\omega)\right)}\left(\sin\left(\theta/2\right)\right)^{1/\omega},$ (6.4)
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\begin{cases}\frac{1}{2\pi}\,_{2}F_{1}\left(\frac{1}{2},\frac{1}{2\omega};\frac{2\omega+1}{2\omega};\frac{t^{2\omega}}{\sin^{2}(\theta/2)}\right),&\quad\mbox{ if }\sin(\theta/2)>t^{\omega},\\[6.0pt] \frac{1}{4\sqrt{\pi}\omega t}\frac{\Gamma\left(1/(2\omega)\right)}{\Gamma\left((\omega+1)/(2\omega)\right)}\left(\sin\left(\theta/2\right)\right)^{1/\omega},&\quad\mbox{ if }\sin(\theta/2)\leq t^{\omega}.\end{cases}$ (6.5)
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left(e^{{\rm i}\theta}\right)=\begin{cases}\frac{1}{2\pi t}\sin\left(\frac{\theta}{2}\right)\arcsin\left(\frac{t}{\sin(\theta/2)}\right),&\quad\mbox{ if }\sin(\theta/2)>t,\\[6.0pt] \frac{1}{4t}\sin\left(\frac{\theta}{2}\right),&\quad\mbox{ if }\sin(\theta/2)\leq t^{.}\end{cases}$ (6.6)

### v2

@@ 4.3. Asymptotically periodic variable Verblunsky coefficients
Here we establish a generalization of Theorem 10 to asymptotically periodic setting. Both approaches presented earlier, the one based on the limiting logarithmic potential and the moment method, extend naturally to this framework.
@@
$\sigma_{t}=\frac{1}{t}\int_{0}^{t}\nu_{\Delta_{s}}{\rm d}s.$
@@
$\mathop{\mathrm{supp}}\nolimits\nu_{\Delta}=\left\{{\rm e}^{{\rm i}\theta}\mid\theta\in[\theta^{(+)},\theta^{(-)}]\cup[-\theta^{(-)},-\theta^{(+)}]\right\},$
@@
$\frac{{\rm d}\nu_{\Delta}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\frac{1}{2\pi}\frac{|\sin\theta|}{\sqrt{\left(\cos\theta^{(+)}-\cos\theta\right)\left(\cos\theta-\cos\theta^{(-)}\right)}},$
for $\pm\theta\in[\theta^{(+)},\theta^{(-)}]$; see [simon_opuc2, Ex. 11.1.5]. Then similarly as in Remark 11, we can deduce the 2-periodic generalization of the formula for the density (4.4) of the average measure (4.21) at the point ${\rm e}^{{\rm i}\theta}$ which reads
@@ 5. Asymptotic zero density for variable OPUC
In this section, we extend the ideas of Theorem 10 to the setting of zeros of OPUC. Recall from Section 2.8 that, even in the non-varying case, the behavior of OPUC zeros is more intricate than that of POPUC. In particular, when $\alpha_{n}\to 0$, the rate of convergence becomes important. More concretely, we prove a variable generalization of the fact that the asymptotic distribution of zeros of OPUC is given by the normalized Lebesgue measure on $\partial\mathbb{D}$} if $\alpha_{n}\to 0$ with $|\alpha_{n}|^{1/n}\to 1$. On the other hand, if the latter condition is altered to $|\alpha_{n}|^{1/n}\to A\in[0,1)$ as in the Mhaskar–Saff theorem, we do not have a variable generalization. This remains an open problem (see Remark 22 below).
@@ 5.1. Balayage
$\int_{\overline{{\mathbb{D}}}}z^{k}{\rm d}\mu(z)=\int_{0}^{2\pi}{\rm e}^{ik\theta}{\rm d}\mathcal{P}(\mu)({\rm e}^{{\rm i}\theta}),\quad k\in\mathbb{N}_{0}.$ (5.1)
@@
$\int_{{\partial\mathbb{D}}}f(z){\rm d}\mathcal{P}(\mu)(z)=\int_{\overline{{\mathbb{D}}}}\mathcal{P}_{*}f(z){\rm d}\mu(z)$
for all continuous functions $f$ on $\partial\mathbb{D}$}, where
$\mathcal{P}_{*}f(r{\rm e}^{{\rm i}\theta})=\begin{cases}\displaystyle\frac{1}{2\pi}\int_{0}^{2\pi}P_{r}(\theta-t)f({\rm e}^{{\rm i}t}){\rm d}t,&\;\mbox{ if }r<1,\\[12.0pt] f({\rm e}^{{\rm i}\theta}),&\;\mbox{ if }r=1,\end{cases}$
@@
$P_{r}(t)=\sum_{n=-\infty}^{\infty}r^{|n|}{\rm e}^{{\rm i}nt}=\frac{1-r^{2}}{1-2r\cos t+r^{2}}$
@@
Recall that $\nu_{n,N}$ denotes the zero-counting measure of $\Phi_{n,N}$, see (1.3).
@@
Since a negative moment of a positive measure supported on $\partial\mathbb{D}$} is the complex-conjugate of the respective positive moment and Laurent polynomials are dense in $C({\partial\mathbb{D}})$ by the Stone–Weierstrass theorem, it is sufficient to show the convergence of the positive moments of the measures in (5.2). By (5.1), the $k$-th moment of $\mathcal{P}(\nu_{n,N})$ equals
$\int_{0}^{2\pi}{\rm e}^{{\rm i}k\theta}{\rm d}\mathcal{P}(\nu_{n,N})\big({\rm e}^{{\rm i}\theta}\big)=\int_{\overline{{\mathbb{D}}}}z^{k}{\rm d}\nu_{n,N}(z)=\frac{1}{n}\mathop{\mathrm{Tr}}\nolimits\left(\mathcal{C}_{n,N}\right)^{k}=\frac{1}{n}\mathop{\mathrm{Tr}}\nolimits\left(\mathcal{C}^{(\beta)}_{n,N}\right)^{k}+o(1),$
@@ 5.2. Asymptotic zero distribution of OPUC with variable Verblunsky coefficients
$\widetilde{\nu}_{n}=\frac{1}{n}\sum_{j=1}^{n}{\delta}_{{\rm e}^{{\rm i}\theta_{j}}},$
where we write $z_{j}=r_{j}{\rm e}^{{\rm i}\theta_{j}}$ for some $r_{j}\geq 0$ and $\theta_{j}\in(-\pi,\pi]$; if $r_{j}=0$ then $\theta_{j}$ can be taken arbitrarily. In the first part of the proof, we show that
$\operatorname*{w-lim\ }_{n\to\infty}(\nu_{n}-\widetilde{\nu}_{n})=0,$ (5.4)
@@
$\operatorname*{w-lim\ }_{n\to\infty}(\widetilde{\nu}_{n}-\mathcal{P}(\nu_{n}))=0.$ (5.5)
Clearly, (5.4) and (5.5) imply the statement.
@@
$\frac{1}{n}\sum_{j=1}^{n}\left|f(z_{j})-f({\rm e}^{{\rm i}\theta_{j}})\right|=\frac{1}{n}\sum_{|z_{j}|\leq 1-\delta}\left|f(z_{j})-f({\rm e}^{{\rm i}\theta_{j}})\right|+\frac{1}{n}\sum_{|z_{j}|>1-\delta}\left|f(z_{j})-f({\rm e}^{{\rm i}\theta_{j}})\right|.$ (5.7)
@@
$\frac{1}{n}\sum_{|z_{j}|\leq 1-\delta}\left|f(z_{j})-f({\rm e}^{{\rm i}\theta_{j}})\right|\leq\frac{2\|f\|_{\overline{{\mathbb{D}}}}}{n}\,\sharp\{j\mid|z_{j}|\leq 1-\delta\},$
@@
$\lim_{n\to\infty}\frac{1}{n}\sum_{|z_{j}|\leq 1-\delta}\left|f(z_{j})-f({\rm e}^{{\rm i}\theta_{j}})\right|=0$
@@
$\frac{1}{n}\sum_{|z_{j}|>1-\delta}\left|f(z_{j})-f({\rm e}^{{\rm i}\theta_{j}})\right|\leq\frac{\epsilon}{n}\,\sharp\{j\mid|z_{j}|>1-\delta\}\leq\epsilon$
@@
$\left|z_{j}-{\rm e}^{{\rm i}\theta_{j}}\right|=1-|z_{j}|<\delta.$
@@
$\lim_{n\to\infty}\frac{1}{n}\sum_{j=1}^{n}\left|f(z_{j})-f({\rm e}^{{\rm i}\theta_{j}})\right|=0,$
@@
2) The verification of (5.5): Both measures $\widetilde{\nu}_{n}$ and $\mathcal{P}(\nu_{n})$ are positive measures supported in $\partial\mathbb{D}$}. Therefore, in order to obtain (5.5), it is sufficient to show that
$\lim_{n\to\infty}\int_{0}^{2\pi}{\rm e}^{{\rm i}kt}{\rm d}\!\left(\widetilde{\nu}_{n}-\mathcal{P}(\nu_{n})\right)({\rm e}^{{\rm i}t})=0$ (5.8)
for all $k\in\mathbb{N}_{0}$. For any $k\in\mathbb{N}_{0}$, one can make use of (5.1) to obtain
$\int_{0}^{2\pi}{\rm e}^{{\rm i}kt}\,{\rm d}\!\left(\widetilde{\nu}_{n}-\mathcal{P}(\nu_{n})\right)({\rm e}^{{\rm i}t})=\int_{0}^{2\pi}{\rm e}^{{\rm i}kt}{\rm d}\widetilde{\nu}_{n}({\rm e}^{{\rm i}t})-\int_{\overline{{\mathbb{D}}}}z^{k}{\rm d}\nu_{n}(z)=\frac{1}{n}\sum_{j=1}^{n}\left({\rm e}^{{\rm i}k\theta_{j}}-z_{j}^{k}\right).$ (5.9)
@@
$\frac{1}{n}\sum_{j=1}^{n}\left|{\rm e}^{{\rm i}k\theta_{j}}-z_{j}^{k}\right|$ $=\frac{1}{n}\sum_{|z_{j}|\leq 1-\delta}\left|{\rm e}^{{\rm i}k\theta_{j}}-z_{j}^{k}\right|+\frac{1}{n}\sum_{|z_{j}|>1-\delta}\left|{\rm e}^{{\rm i}k\theta_{j}}-z_{j}^{k}\right|$
@@
$\operatorname*{w-lim\ }_{n/N\to t}\nu_{n,N}=\frac{1}{t}\int_{0}^{t}\nu_{|\alpha(s)|}{\rm d}s.$
@@
The idea of the proof relies on showing that the property (5.3) holds in the variable setting, i.e the roots $z_{1,n,N},\dots,z_{n,n,N}$ of $\Phi_{n,N}$ are approaching the unit circle up to $o(n)$ exceptions as $n/N\to t$.
@@
$\frac{1}{n_{j}}\log|\alpha_{n_{j}-1,N_{j}}|=\frac{1}{n_{j}}\sum_{k=1}^{n_{j}}\log|z_{k,n_{j},N_{j}}|\leq\frac{\log(1-\delta)}{n_{j}}\,\sharp\{k\mid|z_{k,n_{j},N_{j}}|\leq 1-\delta\}$ (5.12)
@@
It would be very interesting to establish the variable version of the Mhaskar–Saff theorem, that is, to determine the limiting zero density of OPUC when the limit in (5.11) is no longer assumed to be $1$. Addressing this problem likely requires a different approach. We leave this question as an open problem for future.
@@ 6. Examples
As an illustration, we provide examples of POPUC and OPUC with variable Verblunsky coefficients whose asymptotic zero distribution is determined by Theorems 10 and 18 and can be computed in an even more explicit form. This concerns particularly cases when $|\alpha|$ is a monotone function on $(0,t)$.
@@
    If $f$ is strictly increasing in $(0,t)$, then $\sigma_{t}$ is the absolutely continuous measure with $\mathop{\mathrm{supp}}\nolimits\sigma_{t}=\{{\rm e}^{{\rm i}\theta}\mid\theta\in[\theta_{f(0)},2\pi-\theta_{f(0)}]\}$ and
    $\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\int_{0}^{\min(t,f^{-1}(\sin(\theta/2))}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-f^{{2}}(s)}}{\rm d}s,$ (6.1)
@@
    If $f$ is strictly decreasing in $(0,t)$, then $\sigma_{t}$ is the absolutely continuous measure with $\mathop{\mathrm{supp}}\nolimits\sigma_{t}=\{{\rm e}^{{\rm i}\theta}\mid[\theta_{f(t)},2\pi-\theta_{f(t)}]\}$ and
    $\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\frac{1}{2\pi t}\int_{\max(0,f^{-1}(\sin(\theta/2))}^{t}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-f^{{2}}(s)}}{\rm d}s,$ (6.2)
@@
Here $\sigma_{t}$ is the average measure (4.3) and the angle $\theta_{a}$ is defined by (2.17).
@@
Recall that $\mathop{\mathrm{supp}}\nolimits\nu_{f(s)}=\{{\rm e}^{{\rm i}\theta}\mid\theta\in[\theta_{f(s)},2\pi-\theta_{f(s)}]\}$ and therefore, assuming $f$ is increasing, it follows readily from (4.3) that $\mathop{\mathrm{supp}}\nolimits\sigma_{t}=\{{\rm e}^{{\rm i}\theta}\mid\theta\in[\theta_{f(0)},2\pi-\theta_{f(0)}]\}$.
@@
$F_{\sigma_{t}}(\xi)=\sigma_{t}\left(\left\{{\rm e}^{{\rm i}\theta}\mid\theta\in\left[\theta_{f(0)},\xi\right]\right\}\right)=\frac{1}{t}\int_{0}^{t}\nu_{f(s)}\left(\left\{{\rm e}^{{\rm i}\theta}\mid\theta\in\left[\theta_{f(0)},\xi\right]\right\}\right){\rm d}s.$
@@
Suppose $f(s)=s^{\omega}$ and $0<t\leq 1$, where $\omega>0$ is a parameter. This situation corresponds to the Verblunsky coefficients $\alpha_{n,N}=\left(n/N\right)^{\omega}$ for $0\leq n\leq N$, for example. Clearly, $f$ is strictly increasing and, by Proposition 23, $\mathop{\mathrm{supp}}\nolimits\sigma_{t}={\partial\mathbb{D}}$.
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)$ $=\frac{1}{2\pi t}\int_{0}^{t}\frac{\sin\left(\theta/2\right)}{\sqrt{\sin^{2}\left(\theta/2\right)-s^{2\omega}}}{\rm d}s$
$=\frac{1}{2\pi t}(\sin\left(\theta/2\right))^{1/\omega}\int_{0}^{t/(\sin(\theta/2))^{1/\omega}}\frac{{\rm d}u}{\sqrt{1-u^{2\omega}}},$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\frac{1}{2\pi}\,_{2}F_{1}\left(\frac{1}{2},\frac{1}{2\omega};\frac{2\omega+1}{2\omega};\frac{t^{2\omega}}{\sin^{2}(\theta/2)}\right).$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\frac{\left(\sin\left(\theta/2\right)\right)^{\frac{1}{\omega}}}{2\pi t}\,_{2}F_{1}\left(\frac{1}{2},\frac{1}{2\omega};\frac{2\omega+1}{2\omega};1\right)=\frac{\left(\sin\left(\theta/2\right)\right)^{\frac{1}{\omega}}}{2\sqrt{\pi}t}\,\frac{\Gamma\left(\frac{2\omega+1}{2\omega}\right)}{\Gamma\left(\frac{\omega+1}{2\omega}\right)},$
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\begin{cases}\frac{1}{2\pi}\,_{2}F_{1}\left(\frac{1}{2},\frac{1}{2\omega};\frac{2\omega+1}{2\omega};\frac{t^{2\omega}}{\sin^{2}(\theta/2)}\right),&\quad\mbox{ if }\sin(\theta/2)>t^{\omega},\\[8.0pt] \frac{1}{2\sqrt{\pi}t}\frac{\Gamma\left((2\omega+1)/(2\omega)\right)}{\Gamma\left((\omega+1)/(2\omega)\right)}\left(\sin\left(\theta/2\right)\right)^{\frac{1}{\omega}},&\quad\mbox{ if }\sin(\theta/2)\leq t^{\omega}.\end{cases}$ (6.3)
@@
$\frac{{\rm d}\sigma_{t}}{{\rm d}\theta}\left({\rm e}^{{\rm i}\theta}\right)=\begin{cases}\frac{1}{2\pi t}\sin\left(\frac{\theta}{2}\right)\arcsin\left(\frac{t}{\sin(\theta/2)}\right),&\quad\mbox{ if }\sin(\theta/2)>t,\\[6.0pt] \frac{1}{4t}\sin\left(\frac{\theta}{2}\right),&\quad\mbox{ if }\sin(\theta/2)\leq t.\end{cases}$ (6.4)